Hi there, I am trying to solve this problem for a friend. I have been out of school for two long. I know the answer they are looking for is 3 and 3/4 ths. But I don't know how to solve th problem.
if a boy jogs to school at 3 miles per hour and back at the speed of 5 miles per hour and he was gone for 1 hour how many miles did he job round trip

Answer by nerdybill(7384)

if a boy jogs to school at 3 miles per hour and back at the speed of 5 miles per hour and he was gone for 1 hour how many miles did he job round trip
.
Applying the distance formula:
d = rt
.
Let x = time jogging at 3mph
then
1-x = time jogging at 5 mph
.
3x = 5(1-x)
3x = 5-5x
8x = 5
x = 5/8 hour (time jogging at 3mph)
.
time jogging at 5mph:
1-x = 1-5/8 = 8/8-5/8 = 3/8 hour
.
Total distance:
3(5/8) + 5(3/8)
= 15/8 + 15/8
= 30/8
= 15/4
= 3 and 3/4 miles